Five number summary results of diastolic blood pressures (dbp) of a large group of adults are displayed below:

L = = 60, Q1 = 64, Q2 = 75, Q3 = 79, H = 84

What percentage of survey adults have dbp

(i) between 64 and 79? (ii) between 64 and 84? (iii) above 79? (iv) between 75 and 84?
Answer #1

Q1 - 25th perecetile

Q2 - 50 th percentile

Q3 - 75 th percentile

i)

64 is Q1 and 79 is Q3

hence % = 75 -25 = 50 %

ii) 84 is highest

hence % = 100 - 25 = 75 %

iii)

probability = 100- 75 = 25 %

iv)

Q2 = 75

hence probability = 100 -50 = 50 %